OpenStudy (anonymous):

solve the equation 3x-1/ x+2 =7

OpenStudy (anonymous):

3x-1/x+2 = 7 3x-1 = 7(x+2) 3x-1 = 7x+14 -1-14 = 7x -3x -15 = 4x x = -15/4